OpenStudy (anonymous):

Find the matrix of the linear transformation T(f(t)) = f(2t+1) from P2 to P2 with respect to the basis B (script B) = (f1 = 1 +2t^2, f2=1, f3=t)

OpenStudy (anonymous):

halp!

OpenStudy (anonymous):

@No-data

OpenStudy (anonymous):

no-dataaaa!

OpenStudy (anonymous):

It really doesn't

OpenStudy (anonymous):

??

OpenStudy (anonymous):

Fallingangel, don't ever take linear algebra. It is the worstest!

OpenStudy (anonymous):

@no-data see the pdf file.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I think you need to apply the transformation to each vector of the basis.

OpenStudy (anonymous):

and you form a matrix with the coefficients of the result in columns.

OpenStudy (anonymous):

So would the first vector in the basis for f1 be [1,0,2] because of the standard basis [1,x,x^2]?

OpenStudy (anonymous):

I thought that is the way, but I'm checking. if its true.

OpenStudy (anonymous):

Hmm what is the result of apply the transformation to the f_1 vector?

OpenStudy (anonymous):

I got 3 + 8t + 8 t^2

OpenStudy (alexwee123):

ugh linear algebra o.0

OpenStudy (anonymous):

haha true alex. It's sht sh(((ts

OpenStudy (anonymous):

*the

OpenStudy (anonymous):

\[T(f_1(t)) = f_1(2t+1)=1+2(2t+1)^2\]

OpenStudy (anonymous):

I got \(3 + 8t + 8t^2\) am I right?

OpenStudy (anonymous):

sec... I think there's something more you have to do. Let me pull up some pages from a book.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

So it looks like you first have to plug 2t+1 into the x's in your standard P2 polynomial first.

OpenStudy (anonymous):

So you'll have 4t^2+4t+1

OpenStudy (anonymous):

what is your standard polynomial?

OpenStudy (anonymous):

I guess I meant plug in 1 +2t^2 into our polynomial. 4(1+2t^2)^2 +4t +1 = 4(1+2t^2)(1+2t^2) +4t +1 =4(1 + 4t^2 +4t^4) +4t +1

OpenStudy (anonymous):

P2: ax^2 +bx +c

OpenStudy (anonymous):

See I'm really confused. I think I'm going to stop right there b/c I'm not on the right track I don't think.

OpenStudy (anonymous):

You need to remember how to obtain the matrix of a transformation first.

OpenStudy (anonymous):

Always go back to your definitions and well understood theorems brinethery.

OpenStudy (anonymous):

Right now I don't remember well, and I don't have math books at work to help you as I wish. But I think that is the way.

OpenStudy (anonymous):

take some rest if you need it.

OpenStudy (anonymous):

are you allright?

OpenStudy (anonymous):

I'm fine, I just really need help to this question

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

How are you?

OpenStudy (anonymous):

As I said you before, you just need to apply the transformation to each of the vectors of your basis. and form a matrix.

OpenStudy (anonymous):

with the results aligned in columns.

OpenStudy (anonymous):

I mean: \[\left[\begin{matrix} F^T(f_1(t))& F^T(f_2(t)) & F^T(f_3(t)) &\end{matrix}\right]\]

OpenStudy (anonymous):

Sorry I can't see you tube videos at work. =(

OpenStudy (anonymous):

it's blocked.

OpenStudy (anonymous):

That's okay

OpenStudy (anonymous):

but what does it show the video?

OpenStudy (anonymous):

Argentine tango

OpenStudy (anonymous):

My goal in life is to travel to Buenos Aires and learn tango there

OpenStudy (anonymous):

You mean one of the many goals on your life =P

OpenStudy (anonymous):

I have no other goals. If I could just dance then I would be happy lol

OpenStudy (anonymous):

look for punk tango on you tube.

OpenStudy (anonymous):

that is the tango i want to dance haha

OpenStudy (anonymous):

well thank you brinethery.

OpenStudy (anonymous):

I really miss math.. snif snif haha

OpenStudy (anonymous):

see you!

OpenStudy (anonymous):

bye